how many grams of silver iodide, AgI, are produced from 55.3

55.3 g of calcium iodide, CaI2?


1
Expert's answer
2022-05-02T05:08:04-0400

Solution:

The molar mass of CaI2 is 293.887 g/mol

Therefore,

Moles of CaI2 = (55.3 g CaI2) × (1 mol CaI2 / 293.887 g CaI2) = 0.1882 mol CaI2


Balanced chemical equation:

CaI2 + 2AgNO3 → 2AgI + Ca(NO3)2

According to stoichiometry:

1 mol of CaI2 produces 2 mol of AgI

Thus, 0.1882 mol of CaI2 produce:

(0.1882 mol CaI2) × (2 mol AgI / 1 mol CaI2) = 0.3764 mol AgI


The molar mass of AgI is 234.77 g/mol

Therefore,

Mass of AgI = (0.3764 mol AgI) × (234.77 g AgI / 1 mol AgI) = 88.367 g AgI = 88.4 g AgI

Mass of AgI = 88.4 g


Answer: 88.4 grams of silver iodide (AgI)



Short form solution:

Balanced chemical equation:

CaI2 + 2AgNO3 → 2AgI + Ca(NO3)2

According to stoichiometry:

Mass of AgI = (55.3 g CaI2) × (1 mol CaI2 / 293.887 g CaI2) × (2 mol AgI / 1 mol CaI2) × (234.77 g AgI / 1 mol AgI) = 88.35 g AgI = 88.4 g AgI